# Large AI Models for Wireless Physical Layer

### v1

@@ Abstract
Large artificial intelligence models (LAMs) are transforming wireless physical layer technologies through their robust generalization, multitask processing, and multimodal capabilities. This article reviews recent advancements in LAM applications for physical layer communications, addressing limitations of conventional AI-based approaches. LAM applications are classified into two strategies: leveraging pre-trained LAMs and developing native LAMs designed specifically for physical layer tasks. The motivations and key frameworks of these approaches are comprehensively examined through multiple use cases. Both strategies significantly improve performance and adaptability across diverse wireless scenarios. Future research directions, including efficient architectures, interpretability, standardized datasets, and collaboration between large and small models, are proposed to advance LAM-based physical layer solutions for next-generation communication systems.
@@ I Introduction
  2. Obstacle 2: Poor generalization—Practical deployments often face unavoidable mismatches between training and real-world data distributions, degrading NN performance and compromising physical layer reliability.
@@
  4. Obstacle 4: Constrained multimodal processing capabilities—The AI-based physical layer struggles to fully exploit performance gains from other-modal information, particularly sensing data, due to the constrained multimodal processing capabilities.
@@
This article provides a comprehensive review of recent studies on LAMs that focus on the physical layer, covering both the utilization of pre-trained LAMs and the development of native LAMs specifically designed for physical layer. It also seeks to highlight the potential and challenges of physical layer LAMs while providing guidance for future research. The remainder of this article is structured as follows. The next section provides a concise overview of the fundamental concepts of LAMs. Subsequently, we present several application examples of leveraging pre-trained LAMs, including LLMs and LVMs, to enhance physical layer performance. This is followed by a discussion of physical layer-native LAMs, which are trained from scratch using extensive wireless datasets. The final section concludes the article by outlining potential research directions for physical layer LAMs.
@@ II-B Key Advantages of LAMs
Owing to extensive computational resources, advanced algorithms, and large-scale datasets, LAMs significantly outperform smaller AI models, and their primary advantages over smaller models are outlined below.
@@
  4. Advantage 4: Powerful multimodal processing ability—Through the inherent capacity of LAMs’ NNs for high-dimensional representations, massive scaling induces emergent cross-modal grounding, which allows patterns from distinct modalities to interact and translate within a unified processing framework.
@@
BP-LLM [4] Beam pred. GPT-2 Instance norm/patching, embed Introducing Prompt-as-Prefix mechanism 08/2024
@@
2DLAM [10] Delay-Doppler estimation ImageGPT Feature extract, embed $\times$× Introducing LVMs to PHY layer 03/2025
@@
LVM4CSI [14] Channel est., HAR, localization ConvNeXt, DINO-X CSI-to-CV transformation $\times$× $\times$× Proposing LVM-enabled framework for CSI tasks 07/2025
@@
Owing to reference count constraints, only some selected works are included.
III Pre-trained LAMs for Wireless Physical Layer
@@ III-A Motivation
These advantages offer a promising solution to address the challenges of poor generalization immense data dependency, and constrained multimodal processing capabilitie (Obstacles 2, 3 &4). In conventional AI-based physical layer approaches, CSI and other wireless data are often treated as “image” or sequence data, processed by NNs originally developed for CV and NLP. Indeed, these wireless data exhibit high structural similarity to data in CV and NLP. The extensive knowledge acquired by LAMs through vast training datasets can therefore be highly beneficial for physical layer tasks. By inheriting this knowledge, physical layer NNs may require significantly less training data and achieve substantially higher generalization capabilities.
Figure 1: A general framework of pre-trained LAMs for wireless physical layer. This framework adapts pre-trained LAMs with a preprocessing module to align wireless data with the LAM’s input space, and output NN layers to align the output with physical layer tasks. Depending on the task, the LAM’s NN parameters are selectively frozen [14] or fine-tuned [3].
@@ III-B Key Framework
Figure 1 illustrates a general framework of pre-trained LAMs for wireless physical layer. This framework, initially proposed by [3], has been widely adopted by most existing works. This framework consists of three main components, as detailed below.
Preprocessing module: Wireless data, distinct from language and CV modalities, cannot be directly input into pre-trained LAMs and requires preprocessing to align with the LAM’s input format. As most LAMs use a Transformer backbone processing tokenized inputs, the preprocessing module (left part of Figure 1) includes two steps: feature extraction and embedding. Feature extraction uses tailored NN architectures, like stacked FC or convolutional layers in [10], to capture wireless data characteristics. Learnable embedding layers, with attention mechanisms, FC layers, and positional encoding, transform these features into tokens for Transformer-based LAM processing.
Pre-trained LAM: Through training on extensive datasets, pre-trained LAMs acquire broad general knowledge, endowing them with universal modeling capabilities. Consequently, these pre-trained LAMs can serve as the backbone NN module for various physical layer tasks in wireless communications. However, due to significant disparities between the training data used for LAMs and wireless-specific data, directly applying pre-trained LAMs to extract features for physical layer tasks may be suboptimal. As a result, fine-tuning of LAMs is often necessary to enhance performance. Conversely, recent studies suggest that when LVMs or LMMs are employed, the high similarity between natural images and CSI may eliminate the need for fine-tuning, as demonstrated in [14].
@@ III-C1 Pre-trained LLMs for Channel Prediction
The main framework of LLM4CP aligns with the exemplary framework illustrated in Figure 1. To reduce the NN complexity, the LLM4CP is designed to predict future downlink CSI for each antenna pair using its corresponding uplink CSI sequence, rather than the entire CSI. Preprocessing involves normalizing, arranging, and patching frequency- and delay-domain CSI sequences, followed by an NN-based embedding module to generate tokens. These tokens are processed by a pre-trained GPT-2 model, with an FC layer and denormalization to align outputs with the future CSI format. During training, to enhance training efficiency and reducing costs, only specific GPT-2 components, including addition, layer normalization, and positional embedding, are fine-tuned to adapt the LLM for channel prediction, while multi-head attention and feed-forward layers remain frozen to preserve universal knowledge. Across full-sample, few-shot, and generalization tests, simulations confirm LLM4CP’s superior performance against non-AI and other AI methods, at acceptable training and inference costs.
Figure 2: Main framework of the pre-trained LLMs for multimodal sensing-empowered beam prediction, i.e., M2BeamLLM [12], comprising three modules, akin to the general framework illustrated in Figure 1.
@@ III-C2 Pre-trained LVMs for CSI Tasks
CSI is fundamental to the wireless physical layer, with its acquisition and utilization being critical components. Pre-trained LLMs, such as GPT-2, have been employed to improve CSI-related task performance and generalization, leveraging their robust adaptability across diverse tasks [3, 4, 5]. However, the modality gap between CSI and language data necessitates time- and computation-intensive fine-tuning. By visualizing CSI samples in the angular-delay domain, the authors in [14] identifies strong structural similarities with images, proposing LVM4CSI—a pre-trained LVM-based framework that eliminates fine-tuning for CSI-related tasks.
Unlike pre-trained LLM-based approaches that rely on NNs to generate tokens for LLMs, the preprocessing module of LVM4CSI, proposed in [14], transforms CSI data into a CV-compatible image format. For RGB image representation, operations like MATLAB’s jet colormap can directly colorize grayscale CSI samples when only magnitude information is required. Alternatively, a blank channel can be added to a two-channel CSI sample, with real and imaginary parts as separate channels. The RGB CSI “image” is then processed by LVMs (e.g. ConvNeXt) to extract features, followed by a streamlined output module to align these features with the target outcomes. During training, only the NNs in the output module are trained, significantly reducing computational overhead. Simulation results demonstrate that the LVM4CSI framework achieves comparable or superior performance to task-specific NNs.
@@ III-C3 Pre-trained LLMs for Multimodal Sensing-empowered Beam Prediction
Figure 2 depicts the M2BeamLLM framework proposed in [12], comprising three modules, akin to the general framework illustrated in Figure 1. The preprocessing module first encodes sensor data from diverse sources, then fuses and aligns them with the LLM’s input. Similar to LLM4CP, the aligned input is processed by a pre-trained GPT-2 model, with select components frozen and others fine-tuned. The LLM’s output is finally mapped to the beam index. Simulations demonstrate that incorporating sensor data enhances beam prediction performance, with pre-trained LLMs significantly improving accuracy and robustness.
@@ IV-A Motivation
While adapting pre-trained LAMs from domains like CV and NLP offers a pragmatic approach to leverage their powerful capabilities, a fundamental challenge remains: the inherent modality gap between general data (e.g., natural images and text) and wireless signals. Wireless channel data, such as CSI, possesses unique physical properties—e.g., complex-valued numbers, spatial-temporal-frequency correlations, and sparsity in specific domains—that are not present in the datasets used to train LAMs like GPT. This domain gap can render the learned representations from pre-trained LAMs suboptimal for physical layer tasks, as they may not fully capture the underlying physics of signal propagation. Moreover, as existing LLMs and LVMs are designed to address a wide range of tasks with a single model, their complexity is excessive for wireless physical layer applications, rendering them impractical for real-time deployment.
To overcome these limitations, a more direct and potentially more powerful approach has emerged: building native LAMs for the wireless physical layer from the ground up. This approach enables the model to generate universal, rich, and contextualized feature representations (embeddings) of the wireless environment for the utilization of downstream tasks with minimal adaptation. This aligns perfectly with the key advantages of LAMs, particularly robust multi-task processing and excellent few-shot learning (Advantages 2 & 3). Moreover, as LAMs rely minimally on specific NN architectures and incorporate widely adopted blocks like Transformers, they can effectively address the challenges posed by complex NN designs (Obstacle 1).
@@ IV-B Key Framework
Figure 3 depicts a general framework for physical-layer native LAMs. Their development mirrors that of LAMs in NLP and CV, relying on extensive data, robust NNs, and corresponding training as foundational elements.
@@
Robust NNs: Numerous novel NN architectures have been proposed for physical layer tasks. However, research in NLP and CV demonstrates that, with extensive training data, stacked Transformer blocks outperform other meticulously designed NN architectures. Consequently, to develop a robust physical layer LAM, most related studies directly adopt various Transformer blocks, offering a simple yet effective approach.
Task-specific and universal LAMs: As depicted in the left part of Figure 4, physical-layer native LAMs can be designed for specific tasks, including channel prediction, precoding design, and signal detection. Also, leveraging the robust capabilities of LAMs, a single model can effectively perform multiple tasks [9] or extract features for multiple tasks [7, 8].
@@ IV-C Use Cases
The development of native LAMs is proceeding along two distinct strategic paths, each with a different underlying framework. The first path aims to create task-specific models for performance improvement. Conversely, the second path focuses on building highly universal models for broad applicability. The following use cases exemplify these two divergent philosophies.
@@ IV-C1 Task-specific LAM for Wireless Physical Layer
AI-based CSI feedback with an autoencoder framework, studied by 3GPP since 2021, has shown significant potential in reducing feedback overhead and enhancing feedback accuracy. The conventional AI-based feedback NN is trained with CSI data from a certain scenario and evaluated with the data with a similar distribution. This causes a serious generalization problem. Inspired by the powerful generalization ability of LAMs and the mechnasim of AI-based CSI feedback, the authors in [6] propose a prompt-enabled LAM for CSI feedback, which leverages the advantages of both the LAM and the expert knowledge.
The authors in [6] examine the factors contributing to efficient CSI feedback through performance analysis on varied wireless datasets. Their findings indicate that exceptional feedback accuracy derives from AI models’ adeptness at data fitting and their capacity to exploit scenario-specific environmental insights. Leveraging these observations, they propose a pioneering LAM for CSI feedback, incorporating a prompt-based mechanism. This LAM utilizes Transformer architectures and is trained on expansive, multi-scenario datasets. Notably, it embeds expert knowledge—specifically, the average channel magnitude in the angular-delay domain—as a decoder prompt to optimize reconstruction quality. Simulations validate that this prompt-enabled LAM markedly enhances feedback accuracy and adaptability across diverse settings, while minimizing data collection demands in unfamiliar scenarios.
Figure 4: Comparison between task-specific LAMs and multi-task universal LAMs for wireless physical layer applications, demonstrating their respective approaches of employing dedicated AI models for different tasks versus using a single LAM for multiple tasks.

### v2

@@ Abstract
Large artificial intelligence models (LAMs) are transforming wireless physical layer technologies through their robust generalization, multitask processing, and multimodal capabilities. This article reviews recent advancements in applying LAMs to physical layer communications, addressing obstacles of conventional AI-based approaches. LAM-based solutions are classified into two strategies: leveraging pre-trained LAMs and developing native LAMs designed specifically for physical layer tasks. The motivations and key frameworks of these approaches are comprehensively examined through multiple use cases. Both strategies significantly improve performance and adaptability across diverse wireless scenarios. Future research directions, including efficient architectures, interpretability, standardized datasets, and collaboration between large and small models, are proposed to advance LAM-based physical layer solutions for next-generation communication systems.
@@ I Introduction
  2. Obstacle 2: Poor generalization—Practical deployments often face unavoidable mismatches between training and real-world data distributions arising from differences in scenarios, configurations, etc., degrading NN performance and undermining physical layer reliability.
@@
  4. Obstacle 4: Constrained multimodal processing capabilities—The current compact AI-based physical layer struggles to fully exploit performance gains from other-modal information, particularly sensing data, due to the constrained multimodal processing capabilities.
Furthermore, while the computational complexity of NNs remains a significant challenge, the advent of powerful accelerating hardware, combined with model compression methods, enables these models to operate with considerable potential.
@@
This article provides a comprehensive review of recent studies on LAMs that focus on the physical layer, covering both the utilization of pre-trained LAMs and the development of native LAMs specifically designed for physical layer. It also seeks to highlight the potential and challenges of physical layer LAMs while providing guidance for future research. The remainder of this article is structured as follows. The next section provides a concise overview of the fundamental concepts of LAMs and compares pre‑trained and wireless‑native approaches. Subsequently, we present several application examples of leveraging pre-trained LAMs, including LLMs and LVMs, to enhance physical layer performance. This is followed by a discussion of physical layer-native LAMs, which are trained from scratch using extensive wireless datasets. The final section concludes the article by outlining research directions for physical layer LAMs.
Note: Given the focused scope of this paper on LAMs and their advantages over conventional small AI models, we do not provide direct comparisons between AI-based methods and non-AI baselines here. Interested readers are referred to [3] for detailed performance comparisons in this regard.
@@ II-B Key Advantages of LAMs
Owing to extensive computational resources, advanced algorithms, and large-scale datasets, LAMs significantly outperform smaller AI models, and their primary advantages over current smaller models are outlined below.
@@
  4. Advantage 4: Superior multimodal processing ability—Through the inherent capacity of LAMs’ NNs for high-dimensional representations, massive scaling induces emergent cross-modal grounding, which allows patterns from distinct modalities to interact and translate within a unified processing framework.
@@
BP-LLM [9] Beam pred. GPT-2 Instance norm/patching, embed $\times$ Introducing Prompt-as-Prefix mechanism 08/2024
@@
2DLAM [10] Delay-Doppler estimation ImageGPT Feature extract, embed $\times$ Introducing LVMs to PHY layer 03/2025
@@
LVM4CSI [6] Channel est., HAR, localization ConvNeXt, DINO-X CSI-to-CV transformation $\times$ $\times$ Proposing LVM-enabled framework for CSI tasks 07/2025
@@
II-C Comparative Framework: Pre‑trained vs. Wireless-Native
To orient readers before Sections III and IV, this subsection provides a concise framework for choosing between pre-trained LAMs and wireless‑native LAMs. Pre‑trained LAMs adapt general‑purpose models from NLP/CV to physical‑layer tasks via data alignment and task-specific fine‑tuning. Wireless‑native LAMs are trained on wireless signals and scenarios to learn domain-consistent representations.
Selecting between pre-trained and wireless‑native LAMs primarily depends on two factors: modality alignment and computational complexity with real‑time viability. When wireless data such as CSI can be aligned to the source domains of pre‑trained models and latency constraints are relaxed, pre‑trained LAMs enable rapid prototyping and effective multimodal integration with limited labeled data. Under stringent physical-layer latency constraints and device-side power and energy budgets, wireless‑native LAMs, designed with explicit complexity constraints together with model compression and hardware-aware optimization, are preferred. This framing guides the detailed discussions in Sections III and IV. For practitioner assessment of the computational complexity of wireless‑native LAMs, see Section IV-D (Table II); the complexity of pre‑trained backbones is mentioned in Section III-D. The following sections elaborate on these two strategies with representative frameworks and use cases.
III Pre-trained LAMs for Physical Layer
@@ III-A Motivation
These advantages offer a promising solution to address the challenges of poor generalization immense data dependency, and constrained multimodal processing capabilities in conventional/small-scale AI approaches (Obstacles 2, 3 &4). In conventional AI-based physical layer approaches, CSI and other wireless data are often treated as “image” or sequence data, processed by NNs originally developed for CV and NLP. Indeed, these wireless data exhibit high structural similarity to data in CV and NLP. The extensive knowledge acquired by LAMs through vast training datasets can therefore be highly beneficial for physical layer tasks. By inheriting this knowledge, physical layer NNs may require significantly less training data and achieve substantially higher generalization capabilities.
Figure 1: General framework of pre-trained LAMs for physical layer. The framework adapts LAMs using a preprocessing module to align wireless data with input space and output layers to match physical-layer tasks. Depending on the task, LAM parameters are selectively fine-tuned [8] or frozen [6].
@@ III-B Key Framework
Fig. 1 illustrates a general framework of pre-trained LAMs for wireless physical layer. This framework, initially proposed by [8], has been widely adopted by most existing works. This framework consists of three main components, as follows.
Preprocessing module: Wireless data, distinct from language and CV modalities, cannot be directly input into pre-trained LAMs and requires preprocessing to align with the LAM’s input format. As most LAMs use a Transformer backbone processing tokenized inputs, the preprocessing module (left part of Fig. 1) includes two steps: feature extraction and embedding. Feature extraction uses tailored NN architectures, like stacked FC or convolutional layers in [10], to capture wireless data characteristics. Learnable embedding layers, with attention mechanisms, FC layers, and positional encoding, transform these features into tokens for Transformer-based LAM processing.
Pre-trained LAM: Pre-trained LAMs, trained on extensive datasets, acquire broad general knowledge that provides universal modeling capabilities. These models can serve as backbone NN modules for various physical-layer tasks. However, because of the substantial gap between the data used to train LAMs and wireless-specific datasets, directly applying pre-trained LAMs for feature extraction is often suboptimal. In most cases, fine-tuning is necessary to adapt the models and improve performance. Trainable components such as attention layers are commonly adjusted to capture wireless-specific characteristics, as shown in LLM-based approaches [8]. Freezing pre-trained LVMs is considered only when the domain gap is small, for example when CSI is treated as images in LVMs [6]. This approach helps further preserve generalization (Advantage 1), enables few-shot efficiency (Advantage 3), and reduces computational cost, but it remains secondary compared with fine-tuning in most scenarios.
@@ III-C1 Pre-trained LLMs for Channel Prediction
The main framework of LLM4CP is illustrated in Fig. 2. To reduce the NN complexity, the LLM4CP is designed to predict future downlink CSI for each antenna pair using its corresponding uplink CSI sequence, rather than the entire CSI. Preprocessing involves normalizing, arranging, and patching frequency- and delay-domain CSI sequences with attention blocks, followed by an NN-based embedding module to generate tokens. These tokens are processed by a pre-trained GPT-2 model, with an FC layer and denormalization to align outputs with the future CSI format. During training, to enhance training efficiency and reducing costs, only specific GPT-2 components, including addition, layer normalization, and positional embedding, are fine-tuned to adapt the LLM for channel prediction, while multi-head attention and feed-forward layers remain frozen to preserve universal knowledge. Across full-sample, few-shot, and generalization tests, simulations confirm LLM4CP’s superior performance against non-AI and other AI methods, at acceptable training and inference costs.
Figure 3: Main framework of the pre-trained LLMs for multimodal sensing-empowered beam prediction, i.e., M2BeamLLM [13], comprising three modules, akin to the general framework illustrated in Fig. 1.
@@ III-C2 Pre-trained LVMs for CSI Tasks
CSI is fundamental to the wireless physical layer, with its acquisition and utilization being critical components. Pre-trained LLMs, such as GPT-2, have been employed to improve CSI-related task performance and generalization, leveraging their robust adaptability across diverse tasks [8, 4]. However, the modality gap between CSI and language data necessitates time- and computation-intensive fine-tuning. By visualizing CSI samples in the angular-delay domain, the authors in [6] identifies strong structural similarities with images, proposing LVM4CSI—a pre-trained LVM-based framework that eliminates fine-tuning for CSI-related tasks.
Unlike pre-trained LLM-based approaches that rely on NNs to generate tokens for LLMs, the preprocessing module of LVM4CSI, proposed in [6], transforms CSI data into a CV-compatible image format. For RGB image representation, operations like MATLAB’s jet colormap can directly colorize grayscale CSI samples when only magnitude information is required. Alternatively, a blank channel can be added to a two-channel CSI sample, with real and imaginary parts as separate channels. The RGB CSI “image” is then processed by LVMs (e.g. ConvNeXt) to extract features, followed by a streamlined output module to align these features with the target outcomes. During training, only the NNs in the output module are trained, significantly reducing computational overhead. Simulation results demonstrate that the LVM4CSI achieves comparable or superior performance to task-specific NNs.
@@ III-C3 Pre-trained LLMs for Multimodal Sensing-empowered Beam Prediction
Fig. 3 depicts the M2BeamLLM framework proposed in [13], comprising three modules, akin to the general framework illustrated in Fig. 1. The preprocessing module first encodes sensor data from diverse sources, then fuses and aligns them with the LLM’s input. Similar to LLM4CP, the aligned input is processed by a pre-trained GPT-2 model, with select components frozen and others fine-tuned. The LLM’s output is finally mapped to the beam index. Simulations demonstrate that incorporating sensor data enhances beam prediction performance, with pre-trained LLMs significantly improving accuracy and robustness.
III-D Quantitative Complexity Analysis
Pre-trained LAMs (e.g., GPT, LLaMA2, ImageGPT) employed as backbones for wireless feature extraction generally incur substantially higher computational complexity than compact baselines. For example, 2DLAM [10] built on ImageGPT contains 172.56 M parameters [10, Table I], implying significantly greater per-inference computation and latency. Consequently, such models are typically unsuitable for stringent real-time physical-layer tasks, while aligning better with relaxed-latency sensing or offline pipelines.
@@ IV-A Motivation
While adapting pre-trained LAMs from domains like CV and NLP offers a pragmatic approach to leverage their powerful capabilities, a fundamental challenge remains: the inherent modality gap between general data (e.g., natural images and text) and wireless signals. Wireless data, such as CSI, possesses unique physical properties (e.g., complex-valued numbers, spatial-temporal-frequency correlations, and sparsity in specific domains) that are not present in the datasets used to train LAMs. This domain gap renders the learned representations from pre-trained LAMs suboptimal for wireless tasks, as they may not fully capture the underlying wireless physics. Moreover, as existing LLMs and LVMs are designed to address a wide range of tasks with a single model, their complexity is excessive for wireless physical layer applications, rendering them impractical for real-time deployment.
To overcome these limitations, a more direct and potentially more powerful approach has emerged: building native LAMs for the wireless physical layer from the ground up. This approach enables the model to generate universal, rich, and contextualized feature representations (embeddings) of the wireless environment for the utilization of downstream tasks with minimal adaptation. This aligns perfectly with the LAM’s key advantages, particularly robust multi-task processing and excellent few-shot learning (Advantages 2 & 3). Moreover, as LAMs rely minimally on specific NN architectures and incorporate widely adopted Transformer blocks, they can effectively address the challenges of complex NN designs (Obstacle 1).
@@ IV-B Key Framework
Fig. 4 depicts a general framework for physical-layer native LAMs. Their development mirrors that of LAMs in NLP and CV, relying on extensive data, robust NNs, and corresponding training as foundational elements.
@@
Robust NNs: Numerous novel NN architectures have been proposed for physical layer tasks. However, research in NLP and CV demonstrates that, with extensive training data, stacked Transformer blocks outperform other meticulously designed NN architectures. Consequently, to develop a robust physical layer LAM, most related studies directly adopt Transformer blocks, offering a simple yet effective approach.
Task-specific and universal LAMs: As depicted in Fig. 4, physical-layer native LAMs can be designed for specific tasks, including channel prediction, precoding design, and signal detection. Also, leveraging the robust capabilities of LAMs, a single model can effectively perform multiple tasks [14] or extract features for multiple tasks [1, 12].
TABLE II: Performance and complexity comparison of different CSI feedback AI models [5]
Metric Methods
Tiny Simple LAM Prompt+LAM
NMSE (dB) -3.05 -4.59 -6.71 -8.36
Runtime CPU (ms) 3.36 5.20 20.53 21.57
Runtime GPU (ms) 0.90 1.14 2.55 2.64
Parameter number 357,476 563,300 2,546,020 6,705,508
FLOPs (M) 29.17 57.77 334.24 342.56
@@ IV-C Use Cases
The development of native LAMs follows two distinct strategic paths with different frameworks. The first focuses on task-specific models for performance optimization, while the second aims to build universal models for broad applicability. The following use cases illustrate these two approaches.
@@ IV-C1 Task-specific LAM for Wireless Physical Layer
AI-based CSI feedback with an autoencoder framework, studied by 3GPP since 2021, has shown significant potential in reducing feedback overhead and enhancing feedback accuracy. The conventional AI-based feedback NN is trained with CSI data from a certain scenario and evaluated with the data with a similar distribution. This causes a serious generalization problem. Inspired by the powerful generalization ability of LAMs and the mechanism of AI-based CSI feedback, the authors in [5] propose a prompt-enabled LAM for CSI feedback, leveraging the advantages of both the LAMs and expert knowledge.
The authors in [5] examine the factors contributing to efficient CSI feedback through performance analysis on varied wireless datasets. Their findings indicate that exceptional feedback accuracy derives from AI models’ adeptness at data fitting and their capacity to exploit scenario-specific environmental insights. Leveraging these observations, they propose a pioneering LAM for CSI feedback, incorporating a prompt-based mechanism. This LAM utilizes Transformer architectures and is trained on expansive, multi-scenario datasets. Notably, it embeds expert knowledge—specifically, the average channel magnitude in the angular-delay domain—as a decoder prompt to optimize reconstruction quality. At a 144-bit feedback overhead, Table II benchmarks CSI feedback for a prompt‑enabled LAM, its prompt‑free counterpart, and conventional small models. LAMs markedly outperform the small models, with prompts yielding further accuracy gains.